# hyperrefIgnoring empty anchor

### v1

@@ VII Running of the Lensing Anomaly Parameter
respectively, where the parameters $A_{L}$, $B_{L}$, and $n$ are free and $\ell_{0}$ is a pivot multipole number. A similar logarithmic dependence has been studied before in [98] but they were unable to constrain the running parameter $B_{L}$ using then available Planck 2015 temperature anisotropy measurements. We replace the parameter $A_{L}$ defined in CAMB by the expressions in (10),(11). In Fig. (7) we show the fractional change in E-mode power spectrum, $D_{\ell}^{EE}=\ell(\ell+1)\,C_{\ell}^{EE}$ relative to the unlensed spectrum, obtained from the logarithmic and the power-law model for the 4 different values of the running parameter $B_{L}=0,0.1,0.2,0.3$, at the pivot multipole of $\ell_{0}=30$. Given one extra parameter, $n$, in the power-law model the enhancement of power spectrum can be dramatic for bigger values of $n$, in the plot shown, we have, $n=0.5$. We test both of these parametrization first against the P18 dataset alone and then in combination with the Lensing data, and the constraints obtained for the parameters are summarized in Table (5), where we take three different values for the pivot scale, $\ell_{0}=30,100,500$. This choice of pivot scales is motivated by the fact that in previous sections, we observe that deviations in the lensing amplitude comes mostly from $\ell<800$. In addition, for any pivot scale, we obtain two types of constraints; first, we fix $A_{L}=1$ and then we allow $A_{L}$ to freely vary, i.e., we consider a 7- and 8-parameter models, respectively.
@@
It is notable that in our parametrization, it is evident that a change in $l_{0}$ can be absorbed in a redefinition of $A_{L}$ and $B_{L}$. Since both of these parameters are free, the specific selection of $l_{0}$ is, in theory, meaningless. However, due to correlations between the various parameters, in practice, different pivot scale selections may have an impact on the outcomes when both $A_{L}$ and $B_{L}$ are allowed to vary, this fact is effectively demonstrated in Table (5)’s column $\Delta$DIC. It is clear that the choice of $l_{0}$ has only very weak effects on the $\Delta$DIC values for a given model. On the other hand, when $A_{L}$ is fixed at unity, we can no longer claim that the selection of $l_{0}$ is meaningless; in fact, our results are generally affected by this choice.
From Table (5), we see that for the power-law model, $n$ is unconstrained for most cases, and even for the cases where we get constraints on $n$ it comes with large errors, mostly well over 100%. For all four cases considered (i.e., two models applied to two datasets each) when both the parameters $A_{L}$ and $B_{L}$ are allowed to vary the data favors a vanishing $B_{L}$. Upon analyzing the Deviance Information Criterion (DIC), we find weak evidence supporting the power-law model with P18+Lensing. However, we do find strong evidence with P18 for non-zero $B_{L}$ and fixed $A_{L}=1$, both for the power law and logarithmic models. In the most pronounced case, with logarithmic running, $B_{L}$ deviates from zero at $2.6\sigma$, and the DIC gain is $\sim 5$.
@@ VIII Addressing the problematic multipoles
The $\Lambda\text{CDM}+\rm A_{L}$ model when tested against the P18 dataset gives us an anomalous lensing amplitude of $A_{L}=1.183\pm 0.067$, along with a slightly higher $H_{0}$, lower $S_{8}$, and a better DIC than the $\Lambda\text{CDM}$, see Table (6). We remove the low 30 multipoles from the P18 dataset and repeat the analysis to identify constraints on the 7 parameters in order to examine our suspicion that the $-\ell$ measurements of Planck might be the source of this $A_{L}$ anomaly. However, as reported in Table (6), we see that excluding the low multipoles ($\ell<30$) from the P18 dataset results in a doubling of the error bars for $A_{L}$ from 5% to 10%. We observe a similar trend for other parameters as well. Additionally, we find that $\tau$ becomes unconstrained due to the exclusion of lowE dataset (see also [97]) in this analysis. As a result, we discover that it becomes difficult to derive any concrete conclusions from this study by itself, which forces us to find an alternate approach.
@@
To avoid the problem of poor constraints on the parameters, we integrate the reduced P18($\ell<30$) dataset with the WMAP, ACT, and SPT measurements. With this increase in data points, we expect to achieve improved constraints on our parameters. However, before considering the full combination it is crucial to understand how these new datasets and their combinations effect the model parameters of our analysis. When working with the WMAP likelihood  [^1] https://github.com/HTJense/pyWMAP we use a Gaussian prior on $\tau\in\mathcal{N}(0.065,0.015)$. As shown in Table (7), the WMAP data alone cannot constrain the lensing amplitude and only provides an upper limit of $A_{L}<1.16$ within 68% C.L. also we obtain a slightly higher value of $\tau$ as compared to P18 data. Moreover, the WMAP data shows a slight preference for the $\Lambda\text{CDM}$ model over the $\Lambda\text{CDM}+\rm A_{L}$ model, as indicated by the difference in their DIC values. Now, when considering SPT and ACT measurements alongside the WMAP dataset, we observe in Table (7) that all the parameters of the model are nicely constrained. We obtain a lensing amplitude, $A_{L}=0.970^{+0.050}_{-0.058}$, which is consistent with unity within about half a sigma. This combination of the three datasets also shows a preference for the $\Lambda\text{CDM}$ model.
@@
Although the combination of these three datasets provides better-constrained parameters, the error bars are still not as small as those from the full P18 dataset. Hence, we also add the reduced P18($\ell<30$) dataset to our analysis together with the other three CMB measurements and the inferred 68% constrainsts are shown in the last two columns of Table (7). As expected we immediately notice that the constraints obtained are better than the constraints obtained from full P18 alone with the exception of just one parameter $\tau$. This combination gives $A_{L}=1.027\pm 0.042$, consistent with $\Lambda$CDM prediction of unity within $\sim 0.6\sigma$.
At this stage, one might wonder if the reduction in the $A_{L}$ anomaly is due to the inclusion of the WMAP, SPT, and ACT datasets rather than the exclusion of the low multipoles ($\ell<30$) from the P18 dataset. However, this concern is addressed when we examine the constraints obtained from the full P18 dataset along with WMAP, SPT, ACT. As shown in Table 8, the lensing amplitude obtained with the full combined datasets, $A_{L}=1.063\pm 0.035$, is $\sim 2\sigma$ away from unity. Therefore, we can confidently reach to the conclusion that the lensing amplitude anomaly is indeed driven by the lower 30 multipole measurements of P18 dataset.
@@
In Figure (8), we present marginalized 1D posterior for $A_{L}$ obtained from different datasets. Our results indicate that the incorporation of WMAP, ACT, and SPT data substantially mitigates the $A_{L}$ anomaly, rendering it statistically insignificant. Notably, the inclusion of low-z datasets further minimizes the anomaly. With the exception of the full P18 dataset analyzed independently and the case involving the entire dataset combined, the DIC consistently indicates that the $\Lambda$CDM model is preferred over the $\Lambda\text{CDM}+A_{L}$ model for all other datasets and their combinations. In Figure 9 we present likelihood contours for $H_{0}$ and $S_{8}$ for the different combinations of datasets for $\Lambda$CDM model and its extension $\Lambda$CDM$+A_{L}$ model. While these tensions still persist, we can observe some reduction when considering WMAP+ACT+SPT dataset with the $\Lambda$CDM model.
@@ IX Discussion
In the era of precision cosmology theoretical modelling is crucially important; an erroneous determination of one parameter can lead to further erroneous determinations or other – correlated – parameter and thereby to biased scientific conclusions. A reliable inference of $A_{L}\neq 1$ in particular, or of the lensing anomaly in general (i.e. parameterized differently) could possibly imply the presence of an unaccounted-for systematic or more positively, New Physics. In this work we have reanalyzed currently available cosmological data putting special emphasis on the lensing of the CMB by the intervening large scale structure between the last scattering surface and us. The lensing of CMB photons is an important consistency test of the cosmological model, and “delensing” is critical for a reliable detection of primordial gravitational waves via their unique fingerprint in the primordial BB spectrum, e.g. [99].
@@ Acknowledgements
We would like to thank Eleonora Di Valentino for useful correspondence. We acknowledge the Ariel HPC Center at Ariel University for providing computing resources that have contributed to the research results reports reported within this paper. IBD, UK, and AV are supported in part by the “Program of Support of High Energy Physics” Grant by Israeli Council for Higher Education.

### v2

@@ VII Running of the Lensing Anomaly Parameter
respectively, where the parameters $A_{L}$, $B_{L}$, and $n$ are free and $\ell_{0}$ is a pivot multipole number. A similar logarithmic dependence has been studied before in [102] but they were unable to constrain the running parameter $B_{L}$ using then available Planck 2015 temperature anisotropy measurements. We replace the parameter $A_{L}$ defined in CAMB by the expressions in (10), (11). In Fig. 7 we show the fractional change in E-mode power spectrum, $D_{\ell}^{EE}=\ell(\ell+1)\,C_{\ell}^{EE}$ relative to the unlensed spectrum, obtained from the logarithmic and the power-law model for the 4 different values of the running parameter $B_{L}=0,0.1,0.2,0.3$, at the pivot multipole of $\ell_{0}=30$. Given one extra parameter, $n$, in the power-law model the enhancement of power spectrum can be dramatic for bigger values of $n$, in the plot shown, we have, $n=0.5$. We test both of these parametrization first against the P18 dataset alone and then in combination with the Lensing data, and the constraints obtained for the parameters are summarized in Table 5, where we take three different values for the pivot scale, $\ell_{0}=30,100,500$. This choice of pivot scales is motivated by the fact that in previous sections, we observe that deviations in the lensing amplitude come mostly from $\ell<800$. In addition, for any pivot scale, we obtain two types of constraints; first, we fix $A_{L}=1$ and then we allow $A_{L}$ to freely vary, i.e., we consider a 7- and 8-parameter models, respectively.
@@
It is notable that in our parametrization, it is evident that a change in $l_{0}$ can be absorbed in a redefinition of $A_{L}$ and $B_{L}$. Since both of these parameters are free, the specific selection of $l_{0}$ is, in theory, meaningless. However, due to correlations between the various parameters, in practice, different pivot scale selections may have an impact on the outcomes when both $A_{L}$ and $B_{L}$ are allowed to vary, this fact is effectively demonstrated in Table 5’s column $\Delta$DIC. It is clear that the choice of $l_{0}$ has only very weak effects on the $\Delta$DIC values for a given model. On the other hand, when $A_{L}$ is fixed at unity, we can no longer claim that the selection of $l_{0}$ is meaningless; in fact, our results are generally affected by this choice.
From Table 5, we see that for the power-law model, $n$ is unconstrained for most cases, and even for the cases where we get constraints on $n$ it comes with large errors, mostly well over 100%. For all four cases considered (i.e., two models applied to two datasets each) when both the parameters $A_{L}$ and $B_{L}$ are allowed to vary the data favors a vanishing $B_{L}$. Upon analyzing the Deviance Information Criterion (DIC), we find weak evidence supporting the power-law model with P18+Lensing. However, we do find strong evidence with P18 for non-zero $B_{L}$ and fixed $A_{L}=1$, both for the power law and logarithmic models. In the most pronounced case, with logarithmic running, $B_{L}$ deviates from zero at $2.6\sigma$, and the DIC gain is $\sim 5$.
@@ VIII Addressing the problematic multipoles
The $\Lambda\text{CDM}+\rm A_{L}$ model when tested against the P18 dataset gives us an anomalous lensing amplitude of $A_{L}=1.183\pm 0.067$, along with a slightly higher $H_{0}$, lower $S_{8}$, and a better DIC than the $\Lambda\text{CDM}$, see Table 6. We remove the low 30 multipoles from the P18 dataset and repeat the analysis to identify constraints on the 7 parameters in order to examine our suspicion that the $-\ell$ measurements of Planck might be the source of this $A_{L}$ anomaly. However, as reported in Table 6, we see that excluding the low multipoles ($\ell<30$) from the P18 dataset results in a doubling of the error bars for $A_{L}$ from 5% to 10%. We observe a similar trend for other parameters as well. Additionally, we find that $\tau$ becomes unconstrained due to the exclusion of lowE dataset (see also [101]) in this analysis. As a result, we discover that it becomes difficult to derive any concrete conclusions from this study by itself, which forces us to find an alternate approach.
@@
To avoid the problem of poor constraints on the parameters, we integrate the reduced P18($\ell>30$) dataset with the WMAP, ACT, and SPT measurements. With this increase in data points, we expect to achieve improved constraints on our parameters. However, before considering the full combination it is crucial to understand how these new datasets and their combinations effect the model parameters of our analysis. When working with the WMAP likelihood  [^1] https://github.com/HTJense/pyWMAP we use a Gaussian prior on $\tau\in\mathcal{N}(0.065,0.015)$.
As shown in Table 7, the WMAP data alone cannot constrain the lensing amplitude and only provide an upper limit of $A_{L}<1.16$ within 68% C.L. Additionally, we obtain a slightly higher value of $\tau$ compared to the P18 data. Moreover, the WMAP data shows a slight preference for the $\Lambda\text{CDM}$ model over the $\Lambda\text{CDM}+\rm A_{L}$ model, as indicated by the difference in their DIC values. Now, when considering SPT and ACT measurements alongside the WMAP dataset, we observe in Table 7 that all the parameters of the model are nicely constrained. We obtain a lensing amplitude, $A_{L}=0.970^{+0.050}_{-0.058}$, which is consistent with unity within about half a sigma. This combination of the three datasets also shows a preference for the $\Lambda\text{CDM}$ model.
@@
Although the combination of these three datasets provides better-constrained parameters, the error bars are still not as small as those from the full P18 dataset. Hence, we also add the reduced P18($\ell>30$) dataset to our analysis together with the other three CMB measurements and the inferred 68% constraints are shown in the last two columns of Table 7. As expected we immediately notice that the constraints obtained are better than the constraints obtained from full P18 alone with the exception of just one parameter $\tau$. This combination gives $A_{L}=1.027\pm 0.042$, consistent with $\Lambda$CDM prediction of unity within $\sim 0.6\sigma$.
At this stage, one might wonder if the reduction in the $A_{L}$ anomaly is due to the inclusion of the WMAP, SPT, and ACT datasets rather than the exclusion of the low multipoles ($\ell<30$) from the P18 dataset. However, this concern is addressed when we examine the constraints obtained from the full P18 dataset along with WMAP, SPT, ACT. As shown in Table 8, the lensing amplitude obtained with the full combined datasets, $A_{L}=1.063\pm 0.035$, is $\sim 2\sigma$ away from unity. Therefore, we can confidently reach to the conclusion that the lensing amplitude anomaly is indeed driven by the lower 30 multipoles measurements of P18 dataset which is in agreement with previous findings of Planck collaboration [74].
@@
[…]
@@ IX Discussion
In the era of precision cosmology theoretical modeling is crucially important; an erroneous determination of one parameter can lead to further erroneous determinations or other – correlated – parameter and thereby to biased scientific conclusions. A reliable inference of $A_{L}\neq 1$ in particular, or of the lensing anomaly in general (i.e. parameterized differently) could possibly imply the presence of an unaccounted-for systematic or more positively, New Physics. In this work we have reanalyzed currently available cosmological data putting special emphasis on the lensing of the CMB by the intervening large scale structure between the last scattering surface and us. The lensing of CMB photons is an important consistency test of the cosmological model, and “delensing” is critical for a reliable detection of primordial gravitational waves via their unique fingerprint in the primordial BB spectrum, e.g. [103].
@@ Acknowledgements
We would like to thank Eleonora Di Valentino for useful correspondence. This research is partially funded by Ariel University and Afeka Tel-Aviv Academic College of Engineering joint research grant RA2300000410. We acknowledge the Ariel HPC Center at Ariel University for providing computing resources that have contributed to the research results reported within this paper. IBD, UK, and AV are supported in part by the “Program of Support of High Energy Physics” Grant by Israeli Council for Higher Education. MS has been supported by a grant from the Joan and Irwin Jacobs donor-advised fund at the JCF (San Diego, CA). We also thank the anonymous reviewer for their valuable comments and suggestions.